# A volume-of-fluid model for biomass particle pyrolysis

### v1

@@ Abstract
Numerical models of biomass particle pyrolysis focus on either the solid particle evolution or on the surrounding gas-phase dynamics, neglecting the coupled interactions between the two. This work addresses this limitation by proposing a single-grid model that fully resolves both phases without relying on sub-grid-scale correlations. The model adopts an Eulerian representation of the two-phase system, using a Volume-Of-Fluid (VOF) method to track the interface between the biomass and the surrounding gas phase. Solid-phase pyrolysis reactions are included, and a novel approach is proposed to capture the coupling between the evolution of biomass porosity and the particle shrinkage. The anisotropic nature of the biomass particle is accounted for in this multidimensional framework. The resulting model demonstrates mass conservation and numerical convergence. Extensive validation with experimental data shows excellent agreement in terms of mass and temperature profiles and correct volatiles trends. Shrinking profiles reveal correct trends, but they also highlight the need for a better fundamental understanding of the evolution of the biomass structure. Overall, the model takes a step forward in aiding the development of sustainable pyrolysis processes. The code and simulation setups, developed within the open-source Basilisk framework, are made publicly available.
@@
{highlights}
Using Volume-of-Fluid to simulate biomass particle pyrolysis
Resolving both the internal phase and the surrounding environment
Arbitrary moving interface and particle shape
Semi-detailed chemistry and interplay between porosity and volume changes
Extensive validation against experimental measurements
@@ 1 Introduction
To overcome the limitations of the above approaches, this work proposes a unified framework which can resolve the evolution of the solid particle, together with the dynamics of the gas phase. This model resolves the boundary layer at the gas-solid interface, and it does not rely on sub-grid-scale correlations for the heat and mass transfer. The particle evolves considering both porosity changes and the volume-shrinking effects. The resulting formulation is general, and it can be applied to conditions where the interface exchange terms are not known or cannot be easily estimated, which was not possible with the previous models.
@@
This model employs a single grid which describes the solid biomass and the gas phase using an Eulerian approach. In particular, the two-phase system is described using a Volume-Of-Fluid (VOF) model [17], which guarantees a sharp and conservative transport of the interface [18]. The VOF methodology has been employed for simulations of several physical phenomena, such as wave breaking [19], pulsed jets [20], droplet coalescence and breakups [21, 22], bubbly flows [23], droplet evaporation [24, 25], and combustion [26]. In this work, we extend the VOF approach to simulate the evolution of a reacting porous particle. This is made possible through several advancements: i) incorporating porosity within the particle; ii) introducing Darcy–Forchheimer terms into the momentum equations to account for porosity effects; iii) modelling chemical reactions in the solid phase; iv) developing an accurate framework to handle porosity variations and particle shrinkage while ensuring mass conservation. This approach eliminates the numerical instabilities and mass conservation issues typically encountered when using separate grids to represent the solid particle and the surrounding gas phase. The governing equations for this model are described in section 2, together with the underlying assumptions and the biomass-gas boundary conditions. The numerical methods for solving these equations are summarised in section 3. After a brief verification in section 4, the model is used to simulate realistic biomass particles (section 5), which are compared with experimental data in terms of mass and volume changes. The model captures the correct experimental trends and explains the distribution of shrinking and porosity changes inside the particle, which are difficult to measure experimentally.
@@ 2 Mathematical Formulation
Figure 1 displays the sharp distinction between the gas phase and the particle phase. The latter is treated as a single pseudo-phase as it contains a mixture of solid matrix and gas phase within the pores. Mathematically, we distinguish between the two phases by introducing a sharp and a smooth indicator. The sharp indicator function $H$ differentiates between the external gas phase and the porous pseudo-phase, and it is defined as:
@@
This definition implies that the porosity has fractional values within the porous particle, while it is identically equal to $1$ in the gas phase environment. Assuming that the control volume is much larger than the pore size, the porosity results in a smooth function, in analogy to the formulation employed by two-fluid models [29]. The evolution of the porosity is affected by the chemical reactions and the velocity of the pseudo-phase, as discussed in the next sections. Given that the calculation of the porosity is computed by volume averaging over a sufficiently large dimension with respect to the pore size, any random cross-section is statistically representative of the entire 3D structure. Therefore, the volumetric porosity is also equal to the surface porosity, which is used in the calculation of fluxes across the domain [30].
@@ 2.1 Governing equations
$\frac{\partial(\rho_{g}\epsilon_{g}\mathbf{u}_{g})}{\partial t}+\nabla\cdot\left(\rho_{g}\epsilon_{g}\mathbf{u}_{g}\mathbf{u}_{g}\right)=-\nabla(\epsilon_{g}\mathbf{T})+\mathbf{F}_{drag}H$ (6)
@@ 3 Numerical solution strategy
A single set of Navier-Stokes equations is solved for the gas velocity in both phases, using a one-field approach [28]. Details on the solution algorithm for these equations can be found in the work by Popinet [40]. This algorithm is modified in this work in order to include the physical phenomena which characterize the porous systems in this work, as described in the following sections. All the face values in the convective terms are discretised using the second-order upwind method from Bell et al. [41]. Interface transport is performed using a directional split approach [42] in the conservative form proposed by Weymouth and Yue [18]. This method calculates the convective fluxes through a geometric reconstruction of the interface. The same fluxes are also used to solve the convective transport of the scalar fields (e.g. temperature, species mass fractions), which are effectively treated as tracers of the VOF fraction [43]. Furthermore, the solution of the temperature and species field is carried out using a two-field approach. This means that two different governing equations are solved on either side of the interface, with exchange contributions between them. This approach limits unwanted numerical diffusion in the other phase, allows for a precise interface solution, and was found to be substantially more effective when dealing with chemical reactions. Precise temperature and species concentrations are essential for accurate reaction rate calculations, as the exponential temperature dependence can magnify even small variations. Exchange terms between the two fields are found by solving the interface balance equations listed in subsection 2.2 and are introduced as an explicit source term during the scalar diffusion step. This procedure is equivalent to setting Neumann boundary conditions for the particle-environment interface, similarly to what is done in previous works [24, 44].
@@ 3.1 Darcy and Forchheimer contributions
The Darcy and Forchheimer terms are directly included within the solution of the Navier-Stokes equations, and they are non-zero only within the pseudo-phase. Given the very low values of the wood permeability vector $\mathbf{K}$ (commonly in the order of $10^{-14}$ m2), it is essential that these terms are treated either fully or partially implicitly, as an explicit formulation would require unfeasibly low time steps for stability. Using an operator-splitting approach and volumetric averaging, Equation 6 can be simplified to:
@@
with $A=\frac{\mu_{g}\epsilon_{g}}{\mathbf{K}}$, $B=\rho_{g}\frac{1.75}{\sqrt{150\epsilon_{g}^{3}}}\frac{\epsilon_{g}}{\sqrt{\mathbf{K}}}$ and $f$ is the VOF volume fraction defined as $f=\frac{1}{V}\int_{V}Hdv$. The contribution of the Darcy and Forchheimer terms is applied after the predictor step in the Navier-Stokes solution algorithm. Consider $\mathbf{v}_{g}^{*}$ the intermediate gas velocity field found after the predictor step and $\mathbf{v}_{g}^{**}$ the new velocity field after this correction step. Equation 15 is discretised in a semi-implicit way, linearising the Forchheimer contribution using the magnitude of $\mathbf{v}_{g}^{*}$. Doing so, integrating between $t$ and $t+\Delta t$, the gas velocity can be updated as:
@@ 3.3 Chemical kinetics and material properties
The model employs a CHEMKIN format [46], managed through the OpenSMOKE++ library [45], to store and read arbitrarily complex kinetic mechanisms. These mechanisms are used to compute formation rates as well as the required thermodynamic and transport properties, which are evaluated at each time step. In this work, the most recent biomass mechanism developed by CRECK, including the updated cellulose submodel from Debiagi et al. [47], is adopted. The complete kinetic scheme, together with the corresponding transport and physical property data, is provided as supplementary material. Because of their strong variability, the solid-phase density ($\rho_{s}$), thermal conductivity ($\boldsymbol{\lambda}_{s}$), and initial porosity ($\epsilon_{g}^{0}$) were determined separately for each case based on experimental measurements. The initial biomass surrogate composition was estimated from the C/H/O ultimate analysis of the investigated sample, following the procedure of Debiagi et al. [48].
@@ 3.4 Calculation of the shrinking velocity
Summing Equation 19 and 20 results back to Equation 18. $Z(\mathbf{x},t)$ is an arbitrary function of space and time, bounded between 0 and $1$, that distributes the reaction term between volume variation and porosity change. $Z=1$ corresponds to a pure shrinking condition, with no porosity change. On the contrary, $Z=0$ indicates a fixed interface, with only a variation in internal porosity. In principle, the functional form of $Z$ should be related to the physics of the shrinking problem. For example, one could develop an expression correlating the solid properties (i.e., temperature, pressure, and composition) with the importance of shrinking compared to the porosity changes. The functional form of $Z$ serves as a closure model required to solve Equation 18 and can also be discontinuous. In this work, various forms of this function are investigated, with a detailed discussion provided in section 4. Furthermore, a general expression is proposed and tested against real experimental data in section 5. Equation 20 can be solved using a potential flow approximation. The velocity $\mathbf{u}_{p}$ can be set equal to the gradient of a potential field $\psi$:
@@ 3.6 Time-stepping algorithm
For completeness, the whole time-stepping algorithm is reported in algorithm 1. After an initialisation step, where allocation of the necessary memory for the various fields is performed, the time loop begins. At each time step, we first compute the chemistry integration to calculate the interface velocity $\mathbf{u}_{p}$. Once this is established, movement of the interface and the associated tracer fields is performed. Recalculation of the physical properties must now be carried out, as some computational cells could have been emptied of the pseudo-phase tracer or, vice versa, could now include a portion of it. After this step, we proceed with the solution of the Navier-Stokes equation to find the gas velocity $\mathbf{v}_{g}$. The gravity effect can also be included as an additional acceleration term during this part of the algorithm. Basilisk also has a strong adaptive mesh refinement algorithm, which can be optionally used to reduce the total number of computational cells and, therefore, the computational time. After logging the key interest variables, the iteration is repeated until the final time is reached.
@@
4 Mass conservation

### v2

# Multiphase modeling of anisotropic biomass particle pyrolysis accounting for particle deformation and coupled gas-phase dynamics

@@ Abstract
Numerical models of biomass particle pyrolysis focus on either the solid particle evolution or on the surrounding gas-phase dynamics, neglecting the coupled interactions between the two. This work addresses this limitation by proposing a single-grid model that fully resolves both phases without relying on sub-grid-scale correlations. The model adopts an Eulerian representation of the two-phase system, using a Volume-Of-Fluid (VOF) method to track the interface between the biomass and the surrounding gas phase. Solid-phase pyrolysis reactions are included, and a novel approach is proposed to capture the coupling between the evolution of biomass porosity and the particle shrinkage, combining different biomass conversion models into one unique framework. The anisotropic nature of the biomass particle is accounted for in this multidimensional framework. The resulting model is independent of the number and shape of the particle, and demonstrates mass conservation and numerical convergence. Extensive validation with experimental data, collected from wood particles in the centimetre scale and operating temperature between 400-700C, shows excellent agreement in terms of mass and temperature profiles and correct volatiles trends. Predicted char yields fall within 2% error range Shrinking profiles reveal correct trends, with a 10% average error in the final particle shape, but they also highlight the need for a better fundamental understanding of the evolution of the biomass structure. Overall, the model takes a step forward in aiding the development of sustainable pyrolysis processes. The code and simulation setups, developed within the open-source Basilisk framework, are made publicly available.
@@ 1 Introduction
To overcome the limitations of the above approaches, this work proposes a unified framework which can resolve the evolution of the solid particle, together with the dynamics of the gas phase. This model resolves the boundary layer at the gas-solid interface and does not rely on sub-grid-scale correlations for heat and mass transfer. The particle evolves considering both porosity changes and the volume-shrinking effects. The resulting formulation is is independent of the particle shape and number, and it can be applied to conditions where the interface exchange terms are not known or cannot be easily estimated, which was not possible with the previous models.
@@
This model employs a single grid to describe both the solid biomass and the gas phase, using an Eulerian approach. In particular, the two-phase system is described using a Volume-Of-Fluid (VOF) model [33], which has been demonstrated to offer a sharp and conservative transport of the interface [69]. The VOF methodology has been employed for simulations of several physical phenomena, such as wave breaking [22], pulsed jets [40], droplet coalescence and breakups [68, 25], bubbly flows [51], droplet evaporation and combustion [15, 14, 13]. Other works have used VOF-like indicator functions to track evolving fluid-solid interfaces. For example, Maes et al. [45] applied a ”Volume-of-Solid” (VOS) approach at the pore scale within a micro-continuum framework, where the volume fraction field captures the evolving solid-fluid boundary during mineral dissolution. In our work, the VOF indicator tracks the interface between the external gas and a porous pseudo-phase with significant internal flow. At the same time, a separate porosity field describes the internal structure. In this work, the term ”Volume-of-Fluid” is preferred as all numerical methods employed derive directly from the established VOF literature. Furthermore, the emphasis is put on the larger particle scale, extending the VOF approach to simulate the evolution of a reacting porous particle rather than focusing on the evolution of the individual pores. This is made possible through several advancements: i) incorporating porosity within the particle; ii) introducing Darcy–Forchheimer terms into the momentum equations to account for drag effects; iii) modelling chemical reactions in the solid phase; iv) developing an accurate framework to handle porosity variations and particle shrinkage while ensuring mass conservation. This approach limits the numerical instabilities and mass conservation issues typically encountered when using separate grids to represent the solid particle and the surrounding gas phase. The governing equations for this model are described in section 2, together with the underlying assumptions and the biomass-gas boundary conditions. The numerical methods for solving these equations are summarised in section 3. After a brief verification in section 4, the model is used to simulate realistic biomass particles (section 5), which are compared with experimental data in terms of mass and volume changes. The model captures the main experimental trends and it takes a step forward in the understanding of the distribution of shrinking and porosity changes inside the particle, which are difficult to measure experimentally.
The proposed model is independent of the particle shape and number, and provides an excellent candidate for the discovery of sub-scale correlation useful for reactor scale frameworks, like bubbling bed or auger reactors. In particular, models for larger-scale systems are unable to resolve the interface between the biomass and the environment, as well as the evolution of the physical quantities within and around the particles. This information is usually based on literature sub-grid-scale correlations that must be introduced into large-scale models. However, an accurate description of the system requires models tailored for the system and for the operative conditions being solved. In light of this, the present model allows refining the understanding of the particle behaviour to obtain specific correlations for simulating industrial-scale reactors following a multi-scale strategy.
@@ 2 Mathematical Formulation
Figure 1 displays the sharp distinction between the gas phase and the biomass-particle phase. The latter is treated as a single pseudo-phase as it contains a mixture of solid matrix and gas phase within the pores. Mathematically, we distinguish between the different phases by introducing a sharp and a smooth indicator. The sharp indicator function $H$ differentiates between the external gas phase and the porous pseudo-phase, and it is defined as:
@@
This definition implies that the porosity has fractional values within the porous particle, while it is identically equal to $1$ in the gas phase environment. Assuming that the control volume is much larger than the pore size, the porosity results in a smooth function, in analogy to the formulation employed by two-fluid models [47]. The pseudo-phase is treated as a continuum, where overall phase properties are calculated as an average weighted on the porosity scalar field. The evolution of the porosity is affected by the chemical reactions and the velocity of the pseudo-phase, as discussed in the next sections. Given that the calculation of the porosity is computed by volume averaging over a sufficiently large dimension with respect to the pore size, any random cross-section is statistically representative of the entire 3D structure. Therefore, the volumetric porosity is also equal to the surface porosity, which is used in the calculation of fluxes across the domain [53].
@@ 2.1 Governing equations
$\frac{\partial(\rho_{g}\epsilon_{g}\mathbf{u}_{g})}{\partial t}+\nabla\cdot\left(\rho_{g}\epsilon_{g}\mathbf{u}_{g}\mathbf{u}_{g}\right)=-\nabla\cdot(\epsilon_{g}\mathbf{T})+\mathbf{F}_{drag}H$ (6)
@@ 3 Numerical solution strategy
A single set of Navier-Stokes equations is solved for the gas velocity in both phases, using a one-field approach [61]. The solution algorithm is based on a second-order (in space and time) approximate projection method [54, 9], which was developed for incompressible two-phase flows by Popinet [55] and extended to include volumetric sources. The latter makes the velocity field non-solenoidal, thereby making the convergence rate for the particle expansion/shrinkage first-order accurate. This algorithm is further modified in this work in order to include the physical phenomena which characterise the porous systems in this work, as described in the following sections. All the face values in the convective terms are discretised using the second-order upwind method from Bell et al. [9]. Interface transport is performed using a directional split first-order in time forward-Euler approach [58] in the conservative form proposed by Weymouth and Yue [69]. This method calculates the convective fluxes through a geometric reconstruction of the interface. The same fluxes are also used to solve the convective transport of the scalar fields (e.g. temperature, species mass fractions), which are effectively treated as tracers of the VOF fraction [44]. Furthermore, the solution of the temperature and species field is carried out using a two-field approach. This means that two different governing equations are solved on either side of the interface, with exchange contributions between them. This approach limits unwanted numerical diffusion in the other phase, allows for a precise interface solution, and was found to be substantially more effective when dealing with chemical reactions. Precise temperature and species concentrations are essential for accurate reaction rate calculations, as the exponential temperature dependence can magnify even small variations. Exchange terms between the two fields are found by solving the interface balance equations listed in subsection 2.2 and are introduced as an explicit source term during the scalar diffusion step. This procedure is equivalent to setting Neumann boundary conditions for the particle-environment interface, similarly to what is done in previous works [15, 70].
@@ 3.1 Darcy and Forchheimer contributions
The Darcy and Forchheimer terms are directly included within the solution of the Navier-Stokes equations, and they are non-zero only within the pseudo-phase. Given the very low values of the wood permeability vector $\mathbf{K}$ (commonly in the order of $10^{-14}$ m2), it is essential that these terms are treated either fully or partially implicitly, as an explicit formulation would require unfeasibly low time steps for stability. Using an operator-splitting approach it is possible to isolate the effect of the drag term in Equation 6, separating it from the rest of the terms of the Navier–Stokes equation. Noticing that the $\mathbf{F}_{drag}$ term is multiplied by the indicator function $H$, we can use a volumetric average operation to simplify Equation 6 to :
@@
with $A=\frac{\mu_{g}\epsilon_{g}}{\mathbf{K}}$, $B=\rho_{g}\frac{1.75}{\sqrt{150\epsilon_{g}^{3}}}\frac{\epsilon_{g}}{\sqrt{\mathbf{K}}}$, $f$ is the VOF pseudo-phase volume fraction defined as $f=\frac{1}{V}\int_{V}Hdv$, and $\rho_{g}$ and $\mu_{g}$ are respectively the intrinsic internal gas-phase density and viscosity. In the interfacial cells, $\mathbf{F}_{drag}$ only acts on the biomass portion of these cells. Since it is defined per unit volume of biomass, multiplying by $f$ converts it to a force per unit volume of the entire cell. The contribution of the Darcy and Forchheimer terms is applied after the predictor step in the Navier-Stokes solution algorithm. Consider $\mathbf{v}_{g}^{*}$ the intermediate gas velocity field found after the predictor step and $\mathbf{v}_{g}^{**}$ the new velocity field after this correction step. Equation 15 is discretised in a semi-implicit way, linearising the Forchheimer contribution using the magnitude of $\mathbf{v}_{g}^{*}$. Doing so, integrating between $t$ and $t+\Delta t$, the gas velocity can be updated as:
@@
A validation case for the implementation of this term is presented in subsection 4.1
@@ 3.3 Chemical kinetics and material properties
The model employs a CHEMKIN format [37], managed through the OpenSMOKE++ library [18], to store and read arbitrarily complex kinetic mechanisms. These mechanisms are used to compute formation rates as well as the required thermodynamic and transport properties, which are evaluated at each time step. In this work, the most recent biomass mechanism developed by CRECK, including the updated cellulose submodel from Debiagi et al. [21], is adopted. The complete kinetic scheme, together with the corresponding transport and physical property data, is provided as supplementary material. Because of their strong variability, the solid-phase density ($\rho_{s}$), thermal conductivity ($\boldsymbol{\lambda}_{s}$), and initial porosity ($\epsilon_{g}^{0}$) were determined separately for each case based on experimental measurements. The choice of these parameters significantly affects the model output. Literature correlation or direct measurements are preferred whenever possible to maintain a general formulation. The initial biomass surrogate composition was estimated from the C/H/O ultimate analysis of the investigated sample, following the procedure of Debiagi et al. [20]. Permeability measurements are often not available/performed in the same study where pyrolysis data are collected. Therefore, average reported values for the same wood type were used. Moreover, based on the authors’ experience, permeabilities $\mathbf{K<}1\times 10^{-8}$ showed no significant difference, as this low value already implies complete blockage to the external flow. The minor differences are attributable to the internal pressure reached, which has only a minor effect on the variation of physical properties compared to temperature.
@@ 3.4 Calculation of the shrinking velocity
Summing Equation 19 and 20 results back to Equation 18. $Z(\mathbf{x},t)$ is an arbitrary function of space and time, bounded between 0 and $1$, that distributes the reaction term between volume variation and porosity change. $Z=1$ corresponds to a pure shrinking condition, with no porosity change. On the contrary, $Z=0$ indicates a fixed interface, with only a variation in internal porosity. Given that it is not clear what causes and how the shrinking phenomena occur, the proper expression for this function remains unknown. In principle, the functional form of $Z$ should be related to the physics of the shrinking problem. For example, one could develop an expression correlating the solid properties (i.e., temperature, pressure, and composition) with the importance of shrinking compared to the porosity changes. The functional form of $Z$ serves as a closure model required to solve Equation 18 and can also be discontinuous. In this work, various forms of this function are investigated, with a detailed discussion provided in subsection 4.2. Furthermore, a general expression is proposed and tested against real experimental data in section 5. Equation 20 can be solved using a potential flow approximation. The velocity $\mathbf{u}_{p}$ can be set equal to the gradient of a potential field $\psi$:
@@ 3.6 Time-stepping algorithm
For completeness, the whole time-stepping algorithm is reported in algorithm 1. After an initialisation step, where allocation of the necessary memory for the various fields is performed, the time loop begins.
The time step is computed such that it satisfies a Courant-Friedrich-Lewis condition of 0.5, which guarantees the stability of the directional split transport of the VOF scheme [69]. Simulation with an initial surrounding stationary gas phase would imply an unfeasibly large time step, leading to unstable conditions. Therefore, a maximum time step must be set to ensure stability of the temperature and species equations at the initial time step, when the velocity is zero.
At each time step, we first compute the chemistry integration to calculate the interface velocity $\mathbf{u}_{p}$. Once this is established, movement of the interface and the associated tracer fields is performed. Recalculation of the physical properties must now be carried out, as some computational cells could have been emptied of the pseudo-phase tracer or, vice versa, could now include a portion of it. After this step, we proceed with the solution of the Navier-Stokes equation to find the gas velocity $\mathbf{v}_{g}$. The gravity effect can also be included as an additional acceleration term during this part of the algorithm. Basilisk also has a strong adaptive mesh refinement algorithm, which can be optionally used to reduce the total number of computational cells and, therefore, the computational time. The grid is refined with a quad/octree strategy, using a wavelet-based error estimation to avoid issues related to the numerical gradients calculation. This approach is described by Van Hooft et al. [65]. The target fields for adaptation of the simulations in this work are: temperature, velocity, inert gas mass fraction and the biomass main released product volatile fraction (levoglucosan). The refinement is adjusted so that the interface always remains at the maximum level of refinement, to guarantee the correct calculation of the interfacial gradients for the solution of the species and thermal boundary layer.
After logging the key interest variables, the iteration is repeated until the final time is reached.
@@
4 Numerical tests
4.1 Beavers and Joseph Problem
We first want to test the proposed formulation in a non-reactive case, to verify the correct implementation of the Darcy drag term.
The Beavers and Joseph problem [8] is a well-known example in fluid mechanics that examines steady plane flow with an interface between a pure fluid and a porous region that lies parallel to the flow direction. A schematic drawing of the computational domain is depicted in Figure 2. The upper region is a free-flow region, whereas the lower region is occupied by a porous medium. Fluid is entering from the left only in the free flow region, but can exit from both regions at the right boundary. The proportions of the domain and the operating conditions were selected following Betchen et al. [10]. The inlet velocity $U_{in}$ is such that, when the profile is fully developed, a condition of $Re=1$ is achieved in the free flow region, with a denoted average velocity $U$. Figure 3 reports the comparison between the simulated velocity profile and the results obtained by Betchen et al. [10] for two different permeability values. For both cases, consider a constant porosity $\epsilon=0.7$
inlet2H2Hclosed8H8Houtflow
Figure 2: Computational domain for the Beavers and Joseph Problem
1-10.5-0.500.50.51100.250.250.50.50.750.75111.251.251.51.5Refer to caption
y/H
u/U
1-10.5-0.500.50.51100.250.250.50.50.750.75111.251.251.51.5Refer to caption
y/H
u/U
Figure 3: Velocity profile along the $y$ direction: comparison between the simulated profile (line) and the numerical results of Betchen et al. [10] (squares) (a) $K=1\times 10^{-2}\times H^{2}$, $U_{in}=1.17$ (b) $K=1\times 10^{-3}\times H^{2}$, $U_{in}=1.04$
It can be observed that good agreement is found between the two numerical results, highlighting the capability of the proposed framework to be applied to non-reactive test cases. As the permeability decreases, the flow conditions tend to complete blockage of the fluid. Small discrepancies may be related because Betchen et al. [10] do not explicitly report the actual inlet velocity $U_{in}$ used in their simulation; therefore, this value may differ slightly between the two numerical tests. The results obtained are also in line with the numerical work of Costa et al. [17]. In contrast to the previous works, the framework proposed here allows for a moving interface, possibly enabling the study of the time evolution of the system with a receding porous-medium interface.
4.2 Mass conservation